OpenStudy (anonymous):

f(x,y)=x^2+y and c(t)=e^t,e^-t write the composition f(c(t)) as a function of t, then differentiating with respect to t

OpenStudy (anonymous):

When c(t) = e^t, f(c(t)) = e^(2t) Differentiating e^(2t) w.r.t. t, 2e^(2t) When c(t) = e^-t, f(c(t)) = e^(-2t) Differentiating e^(-2t) w.r.t. t, -2e^(-2t)

OpenStudy (anonymous):

ok i see! Then to compute d/dt f(c(t)) i use the what I differentiated w.r.t. t?

OpenStudy (anonymous):

I have already differentiated with respect to t. f(c(t)) just means that where x is in f(x) you just replace it with the value of c(t). So by doing this you get a function dependent on t, which like you said, you differentiate w.r.t. to t.

OpenStudy (anonymous):

oh now i see what you mean. Thank you!

OpenStudy (anonymous):

Oh i'm really really sorry. f(c(t)) means where there is x substitute e^t and where there is y substitute e*-t

OpenStudy (anonymous):

e^-t for y.

OpenStudy (anonymous):

oh ok! so I replace the x and the y with the c(t) values and differentiate w.r.t. t?

OpenStudy (anonymous):

so then f(c(t)) becomes e^2t+e(-t) Now differentiate it with respect to t, we get 2e^2t-e^-t

OpenStudy (anonymous):

ok that makes more sense! so if i wanted to compute d/dt f(c(t)) using the chain rule would i use the formula d/dt f(c(t))=df/dx(dx/dt)+df/dy(dy/dt)

OpenStudy (anonymous):

same f(x,y) and C(t)

OpenStudy (anonymous):

yes, exactly. Either way works! :)